OpenStudy (anonymous):

A recipe that makes 6 servings calls for 1/2 cup of butter. Nancy modifies the recipe so that it can serve 8. A. 3/8 B. 2/3 C. 1 1/3 D. 3 D. How many cups of butter does Nancy need?

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

The original recipe has a ratio of 6 servings for a 1/2 cup of butter. If she's modifying the recipe to make 8 servings, the you have to find an equivalent ratio for 8. 6 : 1/2 8 : ___? If you can find out how much you needed to multiply the 6 by to get 8, then all you have to do is multiply the 1/2 by the same number to get the new number of cups.

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

Can you help me do it step-by-step?

OpenStudy (aripotta):

idk, i'd use a proportion personally

OpenStudy (aripotta):

\[\frac{ 6 }{ 0.5}=\frac{ 8 }{ x }\]

OpenStudy (aripotta):

then cross-multiply to get 8(0.5) = 6x and solve for x

OpenStudy (anonymous):

I'm still not understanding it completely

OpenStudy (anonymous):

so i divide 6 and .5?

OpenStudy (aripotta):

\[\frac{ amount\ of\ servings }{ amount\ of\ butter } \; \frac{ 6 }{ 0.5 }=\ \frac{ 8 }{ x }\]

OpenStudy (aripotta):

8(0.5) = 6x 4 = 6x x = 4/6

OpenStudy (aripotta):

and then of course, 4/6 can be simplified to 2/3

OpenStudy (aripotta):

i hope that made sense?

OpenStudy (anonymous):

Yes it did Thank you!

OpenStudy (aripotta):

no problemo :)

OpenStudy (shadowcain):

1/2 ÷ 6 = 1/2 ÷ 6/1 = (1 × 1) / (6 × 2) = 1/12 per serving 1/12 x 8 = 1/12 × 8/1 = (1 × 8) / (12 × 1) = 8/12 = 2/3 To divide fractions you have to flip the second fraction and then multiply. So it will look like this 1/2 * 1/6 and you get 1/12 this is how much you need for 1 serving now multiply that by 8 and you get 8/12 now simplify and you get 2/3 as your answer.